OpenStudy (anonymous):

In the y=3/2x+b what does the b equal

OpenStudy (anonymous):

The b has a y-intercept of (4,5)

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

No that's fine keep going

OpenStudy (anonymous):

Can you put all of that in a simpler way please

OpenStudy (anonymous):

So when I do that do I get b=1

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

Tell me one more time, the y-intercept is (4,5), b=(4,5) And your equation is y=3/2x+b Correct?

OpenStudy (anonymous):

Can you show me the question specifically so I can see what it is asking, if you want.

OpenStudy (anonymous):

I put that wrong, I meant to say does b=-1

OpenStudy (anonymous):

Ok well b=-1 is saying that it has a y-intercept at (0,-1) correct? Follow along with me here.

OpenStudy (anonymous):

Ok

OpenStudy (dan815):

hey i think the question is wrong, it shud be saying that is the point (4,5)

OpenStudy (dan815):

we can find what the 3 intercept must be if the line goes through the point (4,5)

OpenStudy (dan815):

what the B*, or y interecept must be

OpenStudy (dan815):

y=3/2x+b (4,5) plug this point in and see what b has to be 5=3/2(4)+b b=5-6=-1 so y=3/2x-1

OpenStudy (anonymous):

I already know the answer to that already I need help with another ques

OpenStudy (anonymous):

@dan815 can explain this well, thanks for taking over =0) I was getting confused, I'm unsure of the question haha

OpenStudy (anonymous):

The one I just sent u

OpenStudy (anonymous):

@dan815